If you have 500 ml of 28% salt-solution, how many milliliters of it should you mix with water in order to create 250 ml of a 17% salt-solution? I know that there is 140 salt in the 500 ml mixture and we need 42.5 salt in the final 250 ml mixture but I am unsure of how to set up the ratio problem
Answer by rothauserc(4718)

we are given:
If you have 500 ml of 28% salt-solution, how many milliliters of it should you mix with water in order to create 250 ml of a 17% salt-solution?
you correctly stated that
.28 * 500 = 140 and
.17 * 250 = 42.5
so we are looking for the solution to
140/500 = 42.5/x
cross multiply and we get x =151.79 ml

140x = 500*42.5
x = 151.79 ml